OpenStudy (anonymous):

The table below represents a linear function f(x) and the equation represents a function g(x): x f(x) -1 -11 0 -1 1 9 g(x) g(x) = 5x + 1 Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). (6 points) Part B: Which function has a greater y-intercept? Justify your answer. (4 points)

OpenStudy (imstuck):

Can you make your table a little more organized, please? What exactly are the x points and which are their y's?

OpenStudy (anonymous):

OK

OpenStudy (imstuck):

Are the points (-1,-11), (0,-1), (1,9)?

OpenStudy (anonymous):

YES

OpenStudy (imstuck):

Ok, yeah those are the points. You just graph them to see that they are linear. To find the slope of the function using the points in your table, use the slope formula which is

OpenStudy (anonymous):

still confused @IMStuck

OpenStudy (imstuck):

It had to be RIGHT THEN that my computer chose to shut down and update. Sorry. Anyways, the slope formula is:

OpenStudy (imstuck):

\[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]using points from your table. Let's use points (0,-1) and (1,9) to find slope.

OpenStudy (imstuck):

\[m=\frac{ 9-(-1) }{ 1-0 }=\frac{ 10 }{ 1 }=10\]The slope is 10 from the table. Got that?

OpenStudy (anonymous):

The slope is 10 , got it

OpenStudy (anonymous):

is that for part a

OpenStudy (imstuck):

In g(x), the slope is given you in the way in which the equation is written. y = mx + b, where m is the slope. In your equation, y = 5x + 1, your slope is 5. And yes, what we did with the values from the table is for part a.

OpenStudy (anonymous):

are we all done?

OpenStudy (imstuck):

No, because we didn't do part b yet. We are done with part a.

OpenStudy (anonymous):

ok i see now

OpenStudy (imstuck):

For part b, we need to use the point-slope formula to find the equation and then the y-intercept. We need this formula for the point-slope:

OpenStudy (imstuck):

\[y-y _{1}=m(x-x _{1})\]The value of y1 and x 1 come from a point that you choose from the table. Let's use the point (1,9) cuz it doesn't have any negatives in it. You fill in the x1 with 1, the y1 with 9, and the m with your slope of 10, like this:

OpenStudy (anonymous):

ok

OpenStudy (imstuck):

\[y-9=10(x-1)\]\[y-9=10x-10\]\[y=10x-1\]Now this is also in slope-intercept form y = mx + b, where m is your slope (10) and b is your y intercept. Here, the -1 is in the b position, so your y intercept for this equation is -1. In g(x) = 5x+1, the +1 is in the b position, so your y intercept for g(x) is +1.

OpenStudy (imstuck):

g(x) has a greater y intercept because it is positive 1 and the y intercept from your table is -1.

OpenStudy (anonymous):

thanks for ur help , i get it now

OpenStudy (imstuck):

Good! Glad to help!